In a classroom of 30 students, 3 of the students wear wrist watches.

a) If 14 students are selected with replacement, what is the probability that exactly 2 of them wear wrist watches?

b) If 14 students are selected without replacement, what is the probability that exactly 2 of them wear wrist watches?
Answer #1

Ans:

a)with replacement

n=14,p=3/30=0.1

Probability that exactly 2 them wear wrist watches

=14C2*0.1^2*(1-0.1)^12

=0.2570

b)

Probability that exactly 2 them wear wrist watches

=3C2*27C12/30C14

=3*17383860/145422675

=0.3586